# TimeAutoDiff: Combining Autoencoder and Diffusion model for time series tabular data synthesizing

### v1

@@ Abstract
In this paper, we leverage the power of latent diffusion models to generate synthetic time series tabular data. Along with the temporal and feature correlations, the heterogeneous nature of the feature in the table has been one of the main obstacles in time series tabular data modeling. We tackle this problem by combining the ideas of the variational auto-encoder (VAE) and the denoising diffusion probabilistic model (DDPM). Our model named as TimeAutoDiff has several key advantages including (1) Generality: the ability to handle the broad spectrum of time series tabular data from single to multi-sequence datasets; (2) Good fidelity and utility guarantees: numerical experiments on six publicly available datasets demonstrating significant improvements over state-of-the-art models in generating time series tabular data, across four metrics measuring fidelity and utility; (3) Fast sampling speed: entire time series data generation as opposed to the sequential data sampling schemes implemented in the existing diffusion-based models, eventually leading to significant improvements in sampling speed, (4) Entity conditional generation: the first implementation of conditional generation of multi-sequence time series tabular data with heterogenous features in the literature, enabling scenario exploration across multiple scientific and engineering domains. Codes are in preparation for release to the public, but available upon request.
@@ 1 Introduction
Given the importance of synthesizing tabular data, many researchers have put enormous efforts into building tabular synthesizers with high fidelity and utility guarantees. For example, CTGAN [38] and its variants [44, 45] (e.g., CTABGAN, CTABGAN+) have gained popularity for generating tabular data using a Generative Adversarial Networks [4] (GANs). Recently, diffusion-based tabular synthesizers, like Stasy [13], have shown promise, outperforming GAN-based methods in various tasks. Yet, diffusion models [7, 32] were not initially designed for heterogeneous features. New approaches, such as those using Doob’s h-transform [22], TabDDPM [15], and CoDi [17], aim to address this challenge by combining different diffusion models [32, 10] or leveraging contrastive learning [29] to co-evolve models for improved performance on heterogeneous data. Most recently, researchers have used the idea of latent diffusion model, i.e., AutoDiff [34] and TabSyn [42], to model the heterogeneous features in tables, and prove its empirical effectiveness in various tabular generation tasks.
@@ 1.1 Problem Formulation, Challenges, and Contributions
Our goal is to learn the joint distribution of time series tabular data of a $T$-sequence $(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$. Each observation $\mathbf{x}_{j}$, where $\mathbf{x}_{j}:=[\mathbf{x}_{\text{Cont},j},\mathbf{x}_{\text{Disc},j}]$ consists of a $F$-dimensional feature vector that includes both continuous ($\mathbf{x}_{\text{Cont},j}$) and discrete variables ($\mathbf{x}_{\text{Disc},j}$), reflecting the heterogeneous nature of the dataset. Throughout this paper, we assume there are $B$ i.i.d. observed sequences sampled from $\mathbb{P}(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$. We additionally assume that each record in the time-series tabular data includes a timestamp, formatted as ’YEAR-MONTH-DATE-HOURS’. This timestamp serves as an auxiliary variable to aid an inference step in TimeAutoDiff, which will be detailed shortly.
The overview of TimeAutoDiff is provided in Figure 1. Inspired from [34, 42], we take advantage of the latent diffusion model [28]. Our model has two inference stages: (I). training of VAE [14] to project the time series data to latent spaces; (II). training of DDPM [7] to learn the latent distribution of projected time series data. This seemingly simple idea of combining two powerful models resolves several challenges in time series tabular data modeling.
@@
Feature and temporal correlations are important statistical objects that need to be captured when building a time series tabular synthesizer. Yet, due to the heterogeneous nature of tabular data, capturing these correlations is even more demanding than capturing correlations among purely continuous or discrete features. TimeAutoDiff naturally circumvents this challenge by its construction as the diffusion model learns the distribution of latent representations jointly in the continuous space. As long as VAE gives good latent representations of rows in the table, it should be expected that TimeAutoDiff can nicely capture both temporal and feature correlations in the table. Specifically, the encoder in VAE employs $1$-layer transformer [37] to capture feature correlations and Recurrent Neural Network (RNN) [9] to model the temporal dependences of sequences. In Section 4, we empirically verify that TimeAutoDiff indeed captures both correlations well when compared with other State-Of-The-Art (i.e., SOTA) time series tabular synthesizers.
@@
Numerical comparisons of TimeAutoDiff with other models (with publicly available codes), namely, TimeGAN [39], Diffusion-ts [41], TSGM [19], CPAR [43], and DoppelGANger [21] are conducted comprehensively across six real-world datasets under various metrics. Specifically, qualities of generated tables are measured through (1) statistical fidelities to real tables, (2) machine learning utilities in downstream tasks, and (3) sampling times. Specifically, for measuring the fidelities of temporal correlations between synthetic and real heterogeneous timeseries tabular data, we develop a new metric, named Temporal Discriminative Score. Inspired from the paper [39, 43], this metric computes discriminative scores [39] of distributions of inter-row differences [43] in generated and original sequential data. More details will be presented in Section 4.
@@ 1.2 Relevant Literature
GAN-based models. TimeGAN [39] is one of the most popular time series data synthesizers based on GAN framework. Notably, they used the idea of latent GAN employing the auto-encoder for projecting the time series data to latent space and model the distribution of the data in latent space through GAN framework. Recently proposed Electric Health Record (in short EHR)-Safe [40] integrates a GAN with an encoder-decoder module to generate realistic time series and static variables in EHRs with mixed data types, while EHR-M-GAN [18] employs distinct encoders for each data type, enhancing the generation of mixed-type time series in EHRs. Despite these advancements, GAN-based methods still encounter challenges such as non-convergence, mode collapse, generator-discriminator imbalance, and sensitivity to hyperparameter selection, underscoring the need for ongoing refinement in time series data synthesis.
@@ 2 Proposed Model
Variational Auto-Encoder. The pre-processed input data $\mathbf{x}^{\text{Proc}}:=[\mathbf{x}^{\text{Proc}}_{\text{Bin}};\mathbf{x}^{\text{Proc}}_{\text{Cat}};\mathbf{x}^{\text{Proc}}_{\text{Cont}}]\in\mathbb{R}^{B\times T\times F}$ is fed into the VAE. The architecture of the encoder is illustrated in Figure 2. Motivated by TabTransformer [12], we encode the discrete features $\mathbf{x}_{\text{Disc}}:=[\mathbf{x}^{\text{Proc}}_{\text{Bin}};\mathbf{x}^{\text{Proc}}_{\text{Cat}}]=\{\mathbf{x}_{1},\dots,\mathbf{x}_{m}\}$ into a $d$-dimensional (where $d$ is consistently set at $128$ in this paper) continuous representation. This is achieved using a lookup table $\mathbf{e}(\cdot)\in\mathbb{R}^{d}$, referred to as the column feature, with $m$ representing the total number of discrete features.
@@
The goal of introducing embedding for the discrete variables is to allow the model to differentiate the classes in one column from those in the other columns. To embed the continuous features, we employ a frequency-based representation. Let $\nu$ be a scalar value of the $j$-th continuous feature in $\mathbf{x}_{\text{Cont}}^{\text{Proc}}$. Similarly with [23], $\nu$ is projected to the embedding spaces as follows:
@@
Then, FC layers are applied to $\mathbf{x}^{\text{Pre-Out}}$ for each of data types; that is, $\mathbf{x}^{\text{Out}}_{\text{Cont}}:=\text{Sigmoid}(\text{Linear}(\mathbf{x}^{\text{Pre-Out}}))$, $\mathbf{x}^{\text{Out}}_{\text{Bin}}:=\text{Linear}(\mathbf{x}^{\text{Pre-Out}})$, and $\mathbf{x}^{\text{Out}}_{\text{Cat}}:=\text{Linear}(\mathbf{x}^{\text{Pre-Out}})$, denoting continuous, binary, and categorical outputs of the decoder, respectively. Note that for numerical features, sigmoid activation function is taken to match the scale of $\mathbf{x}^{\text{Proc}}_{\text{Cont}}\in[0,1]$. The dimensions of $\mathbf{x}^{\text{Out}}_{\text{Cont}}$ and $\mathbf{x}^{\text{Out}}_{\text{Bin}}$ are the same as their corresponding inputs, but the dimension of $\mathbf{x}^{\text{Out}}_{\text{Cat}}$ is $\sum_{i}K_{i}$ where $K_{i}$ is the number of categories in each categorical variable. We set the dimensions of output features in this way as we used the mean-squared (MSE), binary cross entropy (BCE), and cross-entropy (CE) in Pytorch package. The reconstruction error is defined as:
@@
Similar as in [42], our model does not require the distribution of embeddings $\mathbf{z}^{\text{Lat}}$ to follow a standard normal distribution strictly as the diffusion model additionally handle the distributional modeling in the latent space. Following [42], we adopt the adaptive schedules of $\beta$ with its maximum value set as $0.1$ and minimum as $10^{-5}$, decreasing the $\beta$ by a factor of $0.7$ (i.e., $\beta^{\text{new}}=0.7\beta^{\text{old}}$) from maximum to minimum whenever $\ell_{\text{recons}}$ fails to decrease for a pre-defined number of epochs.
Diffusion for Time Series. TimeAutoDiff is designed to generate the entire time series at once, taking the data of shape $T\times F$ as an input. This should be contrasted to generating rows in the table sequentially (i.e., for instance [20]). We extend the idea of DDPM to make it accommodate the 2D input. For readers’ convenience, we provide an overview of DDPM in Appendix C.
@@
Through (6), cyclic encodings give $16$-dimensional unique representations of timestamps of observed data in the table, and the encoded vector is fed to an MLP block (2) to match the dimension with those of the other inputs’ encodings. The concatenated encodings of ($Z_{n}^{\text{Lat}}$, $\mathbf{n}$, $\mathbf{t}$, ts) are fed into a MLP block, which gives a tensor $\mathbf{N}=[N_{1},N_{2},\dots,N_{T}]^{T}\in\mathbb{R}^{T\times F}$. Inspired from [36], Bi-directional RNN (BRNN) [30] is employed and $\mathbf{N}$ is fed to BRNN as in Figure 2. After the applications of layer-normalization and FC layer, the network $\epsilon_{\theta}$ outputs $[\epsilon_{\theta}(t_{1}),\dots,\epsilon_{\theta}(t_{T})]^{T}\in\mathbb{R}^{T\times F}$ to estimate $\mathbb{E}^{n}$. The sampling process of new latent matrix is deferred to the Appendix D.
@@ 3 Entity Conditional Generation
TimeAutoDiff is capable of accommodating the entity conditional generation in multi-sequence tabular data. This feature is attributed to the unique capability of diffusion model for conditional generations via classifier-free guidance [8] technique. First, the integer labels (i.e., lbl) of the entities (e.g., Stock symbols in nasdaq100 dataset, APPL : 0, MSFT : 1) are encoded to continuous space via the lookup table. After going through the MLP block, the encoded labels are added to the encodings of diffusion step $n$. We denote this label conditioned model as $\mathbf{\epsilon}_{\theta}\big{(}\mathbf{Z}_{n}^{\text{Lat}},n,\mathbf{t},\mathbf{ts},\mathbf{lbl}\big{)}$. Following [8], we jointly train the conditional and unconditional models by randomly setting lbl as the unconditional class identifier $\emptyset$ with some pre-specified probability. (throughout this paper, we set it as $0.1$.) We use a single neural network to parameterize both models. After training, we perform sampling using the following linear combination of the conditional and unconditional noise estimates:
@@
The detailed algorithm for sampling is deferred in Appendix D. For the visual demonstrations of our results, in Figure 4, we plot the marginal distributions of real versus synthetic of card transaction and nasdaq100 datasets. We give the model the labels of the entire entities in each dataset, and the model generates the sequences of corresponding labels. We refer to the model as C-TimeAutoDiff. (Here, the letter ‘C’ stands for conditional.) For numerical comparisons, we evaluate the performance of entity conditional generation using our C-TimeAutoDiff model against baseline models (CPAR, TimeGAN, DoppelGANger, TSGM) and our model TimeAutoDiff through unconditional generation.
@@ 4 Numerical Experiments
In this section, we evaluate our TimeAutoDiff model in two modes: unconditional and conditional generation, in order to verify the quality of generated time series tabular data. To conduct comparative studies, we choose $5$ baseline models: TimeGAN [39], Diffusion-TS [41], TSGM [19], CPAR [43], and DoppelGANger [21]. We use 6 publicly available datasets where Traffic, Pollution, Hurricane, AirQuality are single-sequence data, and card transaction and nasdaq100 are multi-sequence data. Brief introductions of each dataset with the links for the sources are deferred to Appendix A.

### v2

@@ Abstract
In this paper, we leverage the power of latent diffusion models to generate synthetic time series tabular data. Along with the temporal and feature correlations, the heterogeneous nature of the feature in the table has been one of the main obstacles in time series tabular data modeling. We tackle this problem by combining the ideas of the variational auto-encoder (VAE) and the denoising diffusion probabilistic model (DDPM). Our model named as TimeAutoDiff has several key advantages including (1) Generality: the ability to handle the broad spectrum of time series tabular data from single to multi-sequence datasets; (2) Good fidelity and utility guarantees: numerical experiments on six publicly available datasets demonstrating significant improvements over state-of-the-art models in generating time series tabular data, across four metrics measuring fidelity and utility; (3) Fast sampling speed: entire time series data generation as opposed to the sequential data sampling schemes implemented in the existing diffusion-based models, eventually leading to significant improvements in sampling speed, (4) Entity conditional generation: the first implementation of conditional generation of multi-sequence time series tabular data with heterogenous features in the literature, enabling scenario exploration across multiple scientific and engineering domains. Codes for model implementations are available at https://github.com/namjoonsuh/TimeAutoDiff.
@@ 1 Introduction
Given the importance of synthesizing tabular data, many researchers have put enormous efforts into building tabular synthesizers with high fidelity and utility guarantees. For example, CTGAN [38] and its variants [44, 45] (e.g., CTABGAN, CTABGAN+) have gained popularity for generating tabular data using Generative Adversarial Networks [4] (GANs). Recently, diffusion-based tabular synthesizers, like Stasy [13], have shown promise, outperforming GAN-based methods in various tasks. Yet, diffusion models [7, 32] were not initially designed for heterogeneous features. New approaches, such as those using Doob’s h-transform [22], TabDDPM [15], and CoDi [17], aim to address this challenge by combining different diffusion models [32, 10] or leveraging contrastive learning [29] to co-evolve models for improved performance on heterogeneous data. Most recently, researchers have used the idea of a latent diffusion model, i.e., AutoDiff [34] and TabSyn [42], to model the heterogeneous features in tables and prove its empirical effectiveness in various tabular generation tasks.
@@ 1.1 Problem Formulation, Challenges, and Contributions
Our goal is to learn the joint distribution of time series tabular data of a $T$-sequence $(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$. Each observation $\mathbf{x}_{j}$, where $\mathbf{x}_{j}:=[\mathbf{x}_{\text{Cont},j},\mathbf{x}_{\text{Disc},j}]$ consists of an $F$-dimensional feature vector that includes both continuous ($\mathbf{x}_{\text{Cont},j}$) and discrete variables ($\mathbf{x}_{\text{Disc},j}$), reflecting the heterogeneous nature of the dataset. Throughout this paper, we assume there are $B$ i.i.d. observed sequences sampled from $\mathbb{P}(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$. We additionally assume that each record in the time-series tabular data includes a timestamp, formatted as ’YEAR-MONTH-DATE-HOURS’. This timestamp serves as an auxiliary variable to aid an inference step in TimeAutoDiff, which will be detailed shortly.
The overview of TimeAutoDiff is provided in Figure 1. Inspired by [34, 42], we take advantage of the latent diffusion model [28]. Our model has two inference stages: (I). training of VAE [14] to project the time series data to latent spaces; (II). training of DDPM [7] to learn the latent distribution of projected time series data. This seemingly simple idea of combining two powerful models resolves several challenges in time series tabular data modeling.
@@
Feature and temporal correlations are important statistical objects that need to be captured when building a time series tabular synthesizer. Yet, due to the heterogeneous nature of tabular data, capturing these correlations is even more demanding than capturing correlations among purely continuous or discrete features. TimeAutoDiff naturally circumvents this challenge by its construction as the diffusion model learns the distribution of latent representations jointly in the continuous space. As long as VAE gives good latent representations of rows in the table, it should be expected that TimeAutoDiff can nicely capture both temporal and feature correlations in the table. Specifically, the encoder in VAE employs a $1$-layer transformer [37] to capture feature correlations and a Recurrent Neural Network (RNN) [9] to model the temporal dependences of sequences. In Section 4, we empirically verify that TimeAutoDiff indeed captures both correlations well when compared with other State-Of-The-Art (i.e., SOTA) time series tabular synthesizers.
@@
Numerical comparisons of TimeAutoDiff with other models (with publicly available codes), namely, TimeGAN [39], Diffusion-ts [41], TSGM [19], CPAR [43], and DoppelGANger [21] are conducted comprehensively across six real-world datasets under various metrics. Specifically, the qualities of generated tables are measured through (1) statistical fidelities to real tables, (2) machine learning utilities in downstream tasks, and (3) sampling times. Specifically, for measuring the fidelities of temporal correlations between synthetic and real heterogeneous timeseries tabular data, we develop a new metric, named Temporal Discriminative Score. Inspired from the paper [39, 43], this metric computes discriminative scores [39] of distributions of inter-row differences [43] in generated and original sequential data. More details will be presented in Section 4.
@@ 1.2 Relevant Literature
GAN-based models. TimeGAN [39] is one of the most popular time series data synthesizers based on the GAN framework. Notably, they used the idea of latent GAN employing the auto-encoder for projecting the time series data to latent space and model the distribution of the data in latent space through the GAN framework. Recently proposed Electric Health Record (in short EHR)-Safe [40] integrates a GAN with an encoder-decoder module to generate realistic time series and static variables in EHRs with mixed data types, while EHR-M-GAN [18] employs distinct encoders for each data type, enhancing the generation of mixed-type time series in EHRs. Despite these advancements, GAN-based methods still encounter challenges such as non-convergence, mode collapse, generator-discriminator imbalance, and sensitivity to hyperparameter selection, underscoring the need for ongoing refinement in time series data synthesis.
@@ 2 Proposed Model
Variational Auto-Encoder. The pre-processed input data $\mathbf{x}^{\text{Proc}}:=[\mathbf{x}^{\text{Proc}}_{\text{Bin}};\mathbf{x}^{\text{Proc}}_{\text{Cat}};\mathbf{x}^{\text{Proc}}_{\text{Cont}}]\in\mathbb{R}^{B\times T\times F}$ is fed into the VAE. The architecture of the encoder is illustrated in Figure 2. Motivated by TabTransformer [12], we encode the discrete features $\mathbf{x}_{\text{Disc}}:=[\mathbf{x}^{\text{Proc}}_{\text{Bin}};\mathbf{x}^{\text{Proc}}_{\text{Cat}}]=\{\mathbf{x}_{1},\dots,\mathbf{x}_{m}\}$ into a $d$-dimensional (where $d$ is consistently set at $128$ in this paper) continuous representation. This is achieved using a lookup table $\mathbf{e}(\cdot)\in\mathbb{R}^{d}$ with $m$ representing the total number of discrete features.
@@
The goal of introducing embedding for the discrete variables is to allow the model to differentiate the classes in one column from those in the other columns. To embed the continuous features, we employ a frequency-based representation. Let $\nu$ be a scalar value of the $j$-th continuous feature in $\mathbf{x}_{\text{Cont}}^{\text{Proc}}$. Similar to [23], $\nu$ is projected to the embedding spaces as follows:
@@
Then, FC layers are applied to $\mathbf{x}^{\text{Pre-Out}}$ for each of the data types; that is, $\mathbf{x}^{\text{Out}}_{\text{Cont}}:=\text{Sigmoid}(\text{Linear}(\mathbf{x}^{\text{Pre-Out}}))$, $\mathbf{x}^{\text{Out}}_{\text{Bin}}:=\text{Linear}(\mathbf{x}^{\text{Pre-Out}})$, and $\mathbf{x}^{\text{Out}}_{\text{Cat}}:=\text{Linear}(\mathbf{x}^{\text{Pre-Out}})$, denoting continuous, binary, and categorical outputs of the decoder, respectively. Note that for numerical features, the sigmoid activation function is taken to match the scale of $\mathbf{x}^{\text{Proc}}_{\text{Cont}}\in[0,1]$. The dimensions of $\mathbf{x}^{\text{Out}}_{\text{Cont}}$ and $\mathbf{x}^{\text{Out}}_{\text{Bin}}$ are the same as their corresponding inputs, but the dimension of $\mathbf{x}^{\text{Out}}_{\text{Cat}}$ is $\sum_{i}K_{i}$ where $K_{i}$ is the number of categories in each categorical variable. We set the dimensions of output features in this way as we used the mean-squared (MSE), binary cross entropy (BCE), and cross-entropy (CE) in the Pytorch package. The reconstruction error is defined as:
@@
Similar to [42], our model does not require the distribution of embeddings $\mathbf{z}^{\text{Lat}}$ to follow a standard normal distribution strictly, as the diffusion model additionally handles the distributional modeling in the latent space. Following [42], we adopt the adaptive schedules of $\beta$ with its maximum value set as $0.1$ and minimum as $10^{-5}$, decreasing the $\beta$ by a factor of $0.7$ (i.e., $\beta^{\text{new}}=0.7\beta^{\text{old}}$) from maximum to minimum whenever $\ell_{\text{recons}}$ fails to decrease for a predefined number of epochs.
Diffusion for Time Series. TimeAutoDiff is designed to generate the entire time series at once, taking the data of shape $T\times F$ as an input. This should be contrasted to generating rows in the table sequentially (i.e., for instance, [20]). We extend the idea of DDPM to make it accommodate the 2D input. For readers’ convenience, we provide an overview of DDPM in Appendix C.
@@
Through (6), cyclic encodings give $16$-dimensional unique representations of timestamps of observed data in the table, and the encoded vector is fed to an MLP block (2) to match the dimension with those of the other inputs’ encodings. The concatenated encodings of ($Z_{n}^{\text{Lat}}$, $\mathbf{n}$, $\mathbf{t}$, ts) are fed into an MLP block, which gives a tensor $\mathbf{N}=[N_{1},N_{2},\dots,N_{T}]^{T}\in\mathbb{R}^{T\times F}$. Inspired from [36], Bi-directional RNN (BRNN) [30] is employed and $\mathbf{N}$ is fed to BRNN as in Figure 3. After the applications of layer-normalization and FC layer, the network $\epsilon_{\theta}$ outputs $[\epsilon_{\theta}(t_{1}),\dots,\epsilon_{\theta}(t_{T})]^{T}\in\mathbb{R}^{T\times F}$ to estimate $\mathbb{E}^{n}$. The sampling process of the new latent matrix is deferred to Appendix D.
@@ 3 Entity Conditional Generation
TimeAutoDiff is capable of accommodating the entity conditional generation in multi-sequence tabular data. This feature is attributed to the unique capability of the diffusion model for conditional generations via classifier-free guidance [8] technique. First, the integer labels (i.e., lbl) of the entities (e.g., Stock symbols in nasdaq100 dataset, APPL : 0, MSFT : 1) are encoded to continuous space via the lookup table. After going through the MLP block, the encoded labels are added to the encodings of diffusion step $n$. We denote this label-conditioned model as $\mathbf{\epsilon}_{\theta}\big{(}\mathbf{Z}_{n}^{\text{Lat}},n,\mathbf{t},\mathbf{ts},\mathbf{lbl}\big{)}$. Following [8], we jointly train the conditional and unconditional models by randomly setting lbl as the unconditional class identifier $\emptyset$ with some pre-specified probability. (throughout this paper, we set it as $0.1$.) We use a single neural network to parameterize both models. After training, we perform sampling using the following linear combination of the conditional and unconditional noise estimates:
@@
The detailed algorithm for sampling is deferred in Appendix D. For the visual demonstrations of our results, in Figure 4, we plot the marginal distributions of real versus synthetic data of card transaction and nasdaq100 datasets. We give the model the labels of the entire entities in each dataset, and the model generates the sequences of corresponding labels. We refer to the model as C-TimeAutoDiff. (Here, the letter ‘C’ stands for conditional.) For numerical comparisons, we evaluate the performance of entity conditional generation using our C-TimeAutoDiff model against baseline models (CPAR, TimeGAN, DoppelGANger, TSGM) and our model TimeAutoDiff through unconditional generation.
@@ 4 Numerical Experiments
In this section, we evaluate our TimeAutoDiff model in two modes: unconditional and conditional generation, to verify the quality of generated time series tabular data. To conduct comparative studies, we choose $5$ baseline models: TimeGAN [39], Diffusion-TS [41], TSGM [19], CPAR [43], and DoppelGANger [21]. We use 6 publicly available datasets where Traffic, Pollution, Hurricane, and AirQuality are single-sequence data, and card transaction and nasdaq100 are multi-sequence data. Brief introductions of each dataset with the links for the sources are deferred to Appendix A.